OpenStudy (anonymous):

Determine whether each pair of equations represents perpendicular lines y – x = 3 y = –x + 7

OpenStudy (mysesshou):

yes y=x+3 y=-x+7 lines are perpendicular when the other slope is the *negative inverse* of the other. so for first one, slope = m= 1 for the second one, m=-1 so they're perpendicular.

OpenStudy (anonymous):

thanks so much